### v1

@@ 2 Results
The surface-based manipulation in this work involves three strategies: translation, rotation, and flipping (see Fig. 2). To demonstrate the generalizability of surface-based manipulation, we first conducted open-loop teleoperated experiments on a variety of objects and substances of different shapes, sizes, and mechanical properties, including packaged cookies, cakes, a cotton-stuffed toy fish, a roll of tape, and a bag of popcorn (Fig. 3 and Movie S1). We achieved consistent and successful manipulation as our surface-based approach inherently avoids the complexities of grasping, it can accommodate diverse shapes while minimizing stress on the objects.
@@
Building upon the initial demonstrations, we conducted a series of quantitative, vision-based, closed-loop experiments with two modules to evaluate the proposed strategies. Incorporating perception capabilities, our closed-loop control algorithms enhanced the precision and robustness of these manipulation tasks. First, we validated the translation task, an essential foundation for several subsequent tasks, by repositioning objects to desired locations on the entire surface (Fig. 4 and Movie S2). Next, we combined translation and flipping strategies to fully flip planar objects, effectively swapping their top and bottom surfaces, as shown in Fig. 5a-b and Movie S3. The third experiment involves changing the orientation of the object in the middle of two surfaces (Fig. 5c-d and Movie S4). Beyond these pick-and-place tasks, we demonstrated more dexterous manipulations by reshaping a deformable object by folding it to alter its length, as shown in Fig. 6 and Movie S5.
Figure 3: Surface-based manipulation of various types of objects. Rigid objects and the first two white soft objects are manipulated using closed-loop control, while the remaining manipulations use open-loop teleoperation to transition between different states, demonstrating the strategy’s versatility across diverse object types. (a) Rigid object transfer, flipping, and reorientation. (b) Manipulation of soft and deformable objects: adjusting the position or shape of deformable materials in various forms. (c) Manipulation of different objects, from top left to bottom right: a cookie, a cake in a transparent package, a cotton-stuffed toy fish, a roll of tape, and a bag of unevenly distributed popcorn.
@@ 2.1 Translation of Soft and Randomly Shaped Objects
The objective is to move the soft object across the surface, starting from varying initial locations but converging to a consistent reference position by closed-loop control (see Fig. 3b and Fig. 4). In Fig. 4a, we illustrate three sequential manipulation paths, each initiated from a distinct starting position. Upon reaching the target position and achieving a steady state, we manually repositioned the object to a new starting point. Despite the different starting points, each route consistently concluded at the same predetermined location in the top-left quadrant of the surface. As shown Fig. 4b, the object consistently returns to its target after manual displacements, guided by the control strategy and continuous adjustments in surface orientation and height. During Translation 1, positional changes along the X-axis induced significant variations in surface’s pitch. Before Translation 2, placing the object diagonally opposite the target influenced both the surface’s pitch and roll, with roll adjustments more pronounced due to asymmetries in the workspace (see Fig. S7 and Fig. S8). In Translation 3, the primary deviation along the Y-axis mainly affected surface’s roll values, while minor X-axis changes near the end introduced slight pitch alterations. Additionally, height adjustments in each control cycle caused oscillations that moved the object toward its target.
@@
Across six trials (see Movie S2), the object achieved average velocities of 5.62 cm/s along X-axis and 5.55 cm/s along Y-axis, with steady-state errors of 1.14 cm and 0.87 cm, respectively. Subtle positional discrepancies likely arose from differences in achievable tilt angles (Fig. S8). Overall, these results confirm that our manipulation strategy reliably directs objects to their intended targets, even from various starting points, while minimizing deformation of soft objects.
@@ 2.2 Translation and Flipping of Rigid Object
Flipping motions are crucial in both industrial applications such as assembly and food inspection on production lines [56] and domestic robotics for household tasks [57]. Flipping typically requires multiple controllable contact points [58] and necessitates complex manipulators or dual-arm robots working in coordination [59]. It also demands accurate modeling of contact interactions, handling uncertainties, and maintaining grasp stability. Additionally, limitations in the end effector’s workspace and kinematics further complicate the execution of flipping motions. These complexities make traditional manipulation strategies inadequate for efficiently automating flipping tasks. Our approach addresses these issues by removing workspace dependence through nonprehensile dynamic motion enabled by coordinated modular interactions. This strategy executes complex tasks like 180°$\degree$ flipping along a horizontal axis without intricate grasping mechanisms or high DoF.
We tested this approach by first translating an object from a random starting position on Surface 1 to the intersection of two surfaces (see Fig 3a). Knowing the object’s dimensions, we designed the target position so that the object makes contact with Surface 2 while its center of gravity remains on Surface 1. This setup facilitated the flipping maneuver. After the flip, we repositioned the object between the surfaces and repeated these operations to validate both translation and flipping strategies. As shown in Fig. 5a-b and Fig. 3a, the object was initially moved into position, flipped 180° around its Y-axis, and then returned to an intermediate location between the surfaces. Fig. 5b tracks the object’s trajectory, confirming its successful realignment. During the flipping phase, rapid height and orientation adjustments of Surface 2 briefly impacted the object, completing the flip and producing a distinct yaw change around second 8. Overall, these results demonstrate that our surface-based manipulation method efficiently handles complex flipping tasks, expanding the capabilities of surface-based manipulation strategies.
Figure 5: Validation of translation, flipping, and rotation strategies for rigid objects. (a) In translation and flipping experiments, the object (measuring $8cm\times 8cm\times 3cm$ and weighing $110g$) translates from an initial point on Surface 1 to between two surfaces and then flips. After flipping, the object’s position shifts towards Surface 1, and it is translated again between two surfaces. (b), the position of the object in the surface frame is depicted over time. A noticeable $180\degree$ transition in the object’s yaw orientation in the global frame is attributed to the flipping motion. Additionally, changes in the height of Surface 1 and Surface 2 during manipulation are shown.(c) The schematic shows the rotation experiment setup with the object positioned between the two surfaces. The experiment sets four reference angles of $20\degree$, $30\degree$, $60\degree$, and $80\degree$ and aims to rotate the object to the target orientation. (d) The reference angle and the object’s actual orientation are compared. It also highlights the dynamic adjustments in Surface 2’s roll and height during operation.
@@ 2.3 Rotation of Rigid Object
In assembly lines or food production processes, objects often need to be aligned to specific angles before packaging or further handling [60]. To demonstrate the surface-based manipulation approach for reorienting objects, we conducted an experiment to rotate a rigid object within the surfaces around the Z-axis from its initial orientation to various target orientations. Fig. 5c presents a schematic of our experimental setup that utilizes a square prism as the object of manipulation. Building on the success of the transfer strategy validated in the previous experiment, we assumed the object’s initial position to be centered between two surfaces, with its center of gravity resting on one surface, and then adjusted the other surface’s roll and height to achieve a series of target angles. In Fig. 5d, the actual yaw angles of the object and its alignment with the set target angles are shown. A new reference angle is assigned after the object reaches a reference position. Fig. 5d also displays the roll angle and height adjustments of Surface 1. The variation in Surface 1’s roll angle tends to decrease as the discrepancy between the object’s current angle and the reference angle reduces. The maximum surface’s roll angle per cycle, controlled by a low-level PID controller, is capped at $12\degree$. During the experiments, we achieved an average rotational velocity of $1.16^{\degree}/s$. Following each interaction with the object through roll movement, the surface lowers in altitude and adjusts to an offset angle, preparing for the next interaction cycle. Through this experiment, we successfully validated the feasibility of our rotation strategy and demonstrated a novel, nonprehensile approach to reorientation.
@@ 2.4 Shape Manipulation of Deformable Object
Manipulating deformable objects is essential in various fields [61], including manufacturing [62], the food handling [63], healthcare [64], and elderly care [65]. However, automating the handling of deformable objects remains a significant challenge due to the highly nonlinear character of these materials, which makes accurate modeling extremely difficult [21, 66]. Additionally, using grasping-based methods can cause further unpredictable deformations during manipulation [65]. Surface-based manipulation simplifies perception and control requirements because it reduces the complexity associated with modeling contact dynamics and grasp stability. In this experiment, we combined the previously introduced strategies to fold a strip of putty multiple times, relying only on measurements of the object’s length and geometric center, showcasing the endless possibilities of surface-based methods. Extending this strategy to additional or larger modules could potentially enable more complex tasks involving deformable objects, such as folding clothes or kneading dough.
@@
In Fig. 3b and Fig. 6a, we show the process of folding a strip deformable object. In the experiment, we track the length and geometric center of the deformable object strip using a camera and apply a combination of the previously proposed transfer and flipping strategies to achieve the folding of the deformable object. At the initial stage of the experiment, we positioned the object between two surfaces, upon which we implemented our flipping strategy. However, as the object being manipulated is a soft continuum this time, only a part folds rather than undergoing a complete flip. The folding alters the object’s length, and since we are tracking its geometric center, its position coordinates also change instantaneously. The position coordinates of the tracked object, as shown in Fig. 6b, indicate that the first fold occurs at second 1. The variations in object length are displayed in Fig. 6b. Subsequently, we implemented the translation strategy to position the object between the two surfaces for the next folding phase, as shown during the transfer stage in Fig. 6b. Given our knowledge of the object’s length and the aim to fold approximately one-third of it in each attempt, we were able to determine the target reference positions based on the anticipated length of the fold. This approach allowed for precise control of the folding process. We executed the second folding motion once the object reached the reference position. Finally, through the integration of these operations, we successfully transformed a deformable object strip into a rolled-up configuration and changed its length. This demonstrates the versatility of surface-based manipulation and their potential applications across various fields.
@@ 3 Discussion
In this work, we presented a method to overcome the challenge of manipulating soft and deformable objects utilizing modular robotic surfaces and successfully demonstrated three fundamental operations: translation, rotating, and flipping objects on the surface. Unlike traditional grasping methods that require selecting specific grasp points and struggle with large, low-friction, or deformable objects, our surface-based approach inherently extends to objects of different shapes, sizes, and stiffness levels. The hierarchical control structure enables the proposed strategies to be adapted to other hardware setups by using a simple plane as the end-effector and ensuring the hardware has roll, pitch, and height degrees of freedom. For instance, two conventional six-degree-of-freedom robotic arms equipped with plane end-effectors can apply the strategies presented in this study.
Our work offers a new perspective on manipulation, particularly for handling deformable, soft, and fragile objects. This has potential applications in industrial production and the food industry, where packaging and handling of soft, deformable objects like chicken or fish, which are typically slippery and easily damaged, present significant challenges when using traditional gripper-based methods. Using vision-based closed-loop feedback, we achieved robust control by tracking the geometric center position and orientation of objects with known shapes and sizes. However, our current approach faces limitations in more dynamic environments, especially when the shape of the objects is unknown. For example, flipping and rotating may not be effective if the object’s shape is not accurately perceived. Future work could integrate more sophisticated perception algorithms to understand object shapes better and employ learning-based control algorithms to adapt to different objects and more dynamic environments.
@@ 4.1 Surface-Based Integrated Design
The reconfigurable surface implemented in this project comprises two identical robotic modules. Each module has five main components (see Fig. S7(a)). The surface platform is manufactured from an acrylic sheet, serving as the interface for contacting and manipulating objects via reconfigurability. The origami legs consist of a polyimide layer sandwiched between two FR4 layers, adhered together using two adhesive layers activated through a heat press application. The origami design derives from the fundamental waterbomb origami fold that can enable motion across three DoF, similar to the structures and their functionalities in the previous works [35, 53, 54, 55, 67]. The legs facilitate the system’s reconfigurability, with each leg actuated by direct-driven BLDC motors (maxon EC-i 40) linked with 3D-printed PLA connectors. Finally, the base platform fabricated from an acrylic sheet secures the module onto the surface where the modules are fixed. Each motor, allocated to actuate the robotic modules, operates with a $36V$ nominal voltage and can generate a nominal torque of up to $207mN.m$. The STSPIN32F0A module (by STMicroelectronics) performs low-level control of the motors and manages the required current flow from the power supply. A main controller unit (Teensy MicroMod by PJRC and SparkFun) functions to handle high-level control operations and coordinates communication between these two boards. This communication utilizes the MODBUS RS-485 protocol and the Teensy controller’s TTL output. MAX485 module converts the main controller’s TTL output to the MODBUS RS-485 protocol, creating a coherent and reliable communication channel between the low-level actuation control and high-level functions.

### v2

@@ 2 Results
The surface-based manipulation in this work involves three strategies: translation, rotation, and flipping (see Fig. 2). To demonstrate the generalizability of surface-based manipulation, we first conducted open-loop teleoperated experiments on a variety of objects and substances of different shapes, sizes, and mechanical properties, including packaged cookies, cakes, a cotton-stuffed toy fish, a roll of tape, and a bag of popcorn (Fig. 3 and Supplementary Movie 1). We achieved consistent and successful manipulation as our surface-based approach inherently avoids the complexities of grasping, it can accommodate diverse shapes while minimizing stress on the objects.
@@
Building upon the initial demonstrations, we conducted a series of quantitative, vision-based, closed-loop experiments with two modules to evaluate the proposed strategies. Incorporating perception capabilities, our closed-loop control algorithms enhanced the precision and robustness of these manipulation tasks. First, we validated the translation task, an essential foundation for several subsequent tasks, by repositioning objects to desired locations on the entire surface (Fig. 4 and Supplementary Movie 2). Next, we combined translation and flipping strategies to fully flip planar objects, effectively swapping their top and bottom surfaces, as shown in Fig. 5a-b and Supplementary Movie 3. The third experiment involves changing the orientation of the object in the middle of two surfaces (Fig. 5c-d and Supplementary Movie 4). Beyond these pick-and-place tasks, we demonstrated more dexterous manipulations by reshaping a deformable object by folding it to alter its length, as shown in Fig. 6 and Supplementary Movie 5.
Figure 3: Surface-based manipulation of various types of objects. A rigid object (a solid, uniform square prism block) and the first two white soft objects (a malleable slime) are manipulated using closed-loop control, while the remaining objects are manipulated via open-loop teleoperation to transition between different states, demonstrating the strategy’s versatility across diverse object types. (a) Rigid object transfer (T), rotation (R), and flipping (F). (b) Manipulation of soft and deformable objects: adjusting the position or shape of deformable materials in various forms. (c) Manipulation of different objects, from top left to bottom right: a cookie, a cake in a transparent package, a cotton-stuffed toy fish, a roll of tape, and a bag of unevenly distributed popcorn.
@@ 2.1 Translation of Soft and Randomly Shaped Objects
The objective is to move the soft object across the surface, starting from varying initial locations but converging to a consistent reference position by closed-loop control (see Fig. 3b and Fig. 4). In Fig. 4a, we illustrate three sequential manipulation paths, each initiated from a distinct starting position. Upon reaching the target position and achieving a steady state, we manually repositioned the object to a new starting point. Despite the different starting points, each route consistently concluded at the same predetermined location in the top-left quadrant of the surface. As shown Fig. 4b, the object consistently returns to its target after manual displacements, guided by the control strategy and continuous adjustments in surface orientation and height. During Translation 1, positional changes along the X-axis induced significant variations in surface’s pitch. Before Translation 2, placing the object diagonally opposite the target influenced both the surface’s pitch and roll, with roll adjustments more pronounced due to asymmetries in the workspace (see Supplementary Fig. 1 and Supplementary 2). In Translation 3, the primary deviation along the Y-axis mainly affected surface’s roll values, while minor X-axis changes near the end introduced slight pitch alterations. Additionally, height adjustments in each control cycle caused oscillations that moved the object toward its target.
@@
Across six trials (see Supplementary Movie 2), the object achieved average velocities of 5.62 cm/s along X-axis and 5.55 cm/s along Y-axis, with steady-state errors of 1.14 cm and 0.87 cm, respectively. Subtle positional discrepancies likely arose from differences in achievable tilt angles (Supplementary Fig. 2). Overall, these results confirm that our manipulation strategy reliably directs objects to their intended targets, even from various starting points, while minimizing deformation of soft objects.
@@ 2.2 Translation and Flipping of Rigid Object
[…]
We tested this approach using a rectangular cuboid measuring $8\times 8\times 3\ cm^{3}$ and weighing 110 g. The object was first translated from a random starting position on Surface 1 to the intersection of two surfaces (see Fig 3a). Knowing the object’s dimensions, we designed the target position so that the object makes contact with Surface 2 while its center of gravity remains on Surface 1. This setup facilitated the flipping maneuver. After the flip, we repositioned the object between the surfaces and repeated these operations to validate both translation and flipping strategies. As shown in Fig. 5a-b and Fig. 3a, the object was initially moved into position, flipped 180° around its Y-axis, and then returned to an intermediate location between the surfaces. Fig. 5b tracks the object’s trajectory, confirming its successful realignment. During the flipping phase, rapid height and orientation adjustments of Surface 2 briefly impacted the object, completing the flip and producing a distinct yaw change at approximately 8 seconds. Overall, these results demonstrate that our surface-based manipulation method efficiently handles complex flipping tasks, expanding the capabilities of surface-based manipulation strategies.
Figure 5: Validation of translation, flipping, and rotation strategies for rigid objects. (a) In translation and flipping experiments, the object (measuring $8\times 8\times 3\ cm^{3}$ and weighing $110g$) translates from an initial point on Surface 1 to between two surfaces and then flips. After flipping, the object’s position shifts towards Surface 1, and it is translated again between two surfaces. (b), the position of the object in the surface frame is depicted over time. A noticeable $180\degree$ transition in the object’s yaw orientation in the global frame is attributed to the flipping motion. Additionally, changes in the height of Surface 1 and Surface 2 during manipulation are shown.(c) The schematic shows the rotation experiment setup with the object positioned between the two surfaces. The experiment sets four reference angles of $20\degree$, $30\degree$, $60\degree$, and $80\degree$ and aims to rotate the object to the target orientation. (d) The reference angle and the object’s actual orientation are compared. It also highlights the dynamic adjustments in Surface 2’s roll and height during operation.
@@ 2.3 Rotation of Rigid Object
In assembly lines or food production processes, objects often need to be aligned to specific angles before packaging or further handling [chen2023visual]. To demonstrate the surface-based manipulation approach for reorienting objects, we conducted an experiment to rotate a rigid object within the surfaces around the Z-axis from its initial orientation to various target orientations. Fig. 5c presents a schematic of our experimental setup that utilizes a square prism as the object of manipulation. Building on the success of the transfer strategy validated in the previous experiment, we assumed the object’s initial position to be centered between two surfaces, with its center of gravity resting on one surface, and then adjusted the other surface’s roll and height to achieve a series of target angles. In Fig. 5d, the actual yaw angles of the object and its alignment with the set target angles are shown. A new reference angle is assigned after the object reaches a reference position. Fig. 5d also displays the roll angle and height adjustments of Surface 1. The variation in Surface 1’s roll angle tends to decrease as the discrepancy between the object’s current angle and the reference angle reduces. The maximum surface’s roll angle per cycle, controlled by a low-level PID controller, is capped at $12\degree$. During the experiments, we achieved an average rotational velocity of $1.16^{\degree}/s$. Following each interaction with the object through roll movement, the surface lowers in altitude and adjusts to an offset angle, preparing for the next interaction cycle. Through this experiment, we successfully validated the feasibility of our rotation strategy and demonstrated a novel, nonprehensile approach to reorientation.
@@ 2.4 Shape Manipulation of Deformable Object
Manipulating deformable objects is essential in various fields [doi:10.1177/0278364918779698], including manufacturing [NGUYEN2021379], the food handling [satici2022coordinate], healthcare [schmidgall2024general], and elderly care [doi:10.1126/scirobotics.abm6010]. However, automating the handling of deformable objects remains a significant challenge due to the highly nonlinear character of these materials, which makes accurate modeling extremely difficult [doi:10.1126/science.aat8414, yin2021]. Additionally, using grasping-based methods can cause further unpredictable deformations during manipulation [doi:10.1126/scirobotics.abm6010]. Surface-based manipulation simplifies perception and control requirements because it reduces the complexity associated with modeling contact dynamics and grasp stability. In this experiment, we combined the previously introduced strategies to fold a strip of putty multiple times, relying only on measurements of the object’s length and geometric center, showcasing the endless possibilities of surface-based methods. Extending this strategy to additional or larger modules could potentially enable more complex tasks involving deformable objects, such as folding clothes or kneading dough.
@@
In Fig. 3b and Fig. 6a, we show the process of folding a strip deformable object. In the experiment, we track the length and geometric center of the deformable object strip using a camera and apply a combination of the previously proposed transfer and flipping strategies to achieve the folding of the deformable object. At the initial stage of the experiment, we positioned the object between two surfaces, upon which we implemented our flipping strategy. However, as the object being manipulated is a soft continuum this time, only a part folds rather than undergoing a complete flip. The folding alters the object’s length, and since we are tracking its geometric center, its position coordinates also change instantaneously. The position coordinates of the tracked object, as shown in Fig. 6b, indicate that the first fold occurs at second 1. The variations in object length are displayed in Fig. 6b. Subsequently, we implemented the translation strategy to position the object between the two surfaces for the next folding phase, as shown during the transfer stage in Fig. 6b. Given our knowledge of the object’s length and the aim to fold approximately one-third of it in each attempt, we were able to determine the target reference positions based on the anticipated length of the fold. This approach allowed for precise control of the folding process. We executed the second folding motion once the object reached the reference position. Finally, through the integration of these operations, we successfully transformed a deformable object strip into a rolled-up configuration and changed its length. This demonstrates the versatility of surface-based manipulation and their potential applications across various fields. While our method successfully folded a low-stiffness deformable object (Young’s modulus approximately between 100 MPa and 200 MPa [ashraf2022mechanical]), handling more stiff materials presents additional challenges. Increased stiffness can prevent full folding and introduce greater reaction forces at contact points and lead to unintended object displacement. Future adaptations could involve implementing controlled force application, incorporating external constraints to aid bending, or estimating stiffness to refine control strategies.
@@ 3 Discussion
In this work, we presented a method to overcome the challenge of manipulating soft and deformable objects utilizing modular robotic surfaces and successfully demonstrated three fundamental operations: translation, rotating, and flipping objects on the surface. Unlike traditional grasping methods that require selecting specific grasp points and struggle with large, low-friction, or deformable objects, our surface-based approach inherently extends to objects of different shapes, sizes, and stiffness levels. The hierarchical control structure enables the proposed strategies to be adapted to other hardware setups by using a simple plane as the end-effector and ensuring the hardware has roll, pitch, and height DoF. For instance, two conventional 6 DoF robotic arms equipped with plane end-effectors can apply the strategies presented in this study.
Our work presents a new approach to manipulation, particularly for handling deformable, soft, and fragile objects. This has potential applications in industries such as food processing, where packaging and handling items like chicken or fish, which are typically slippery and easily damaged, pose challenges for traditional grasping methods. Using vision-based closed loop feedback, we achieved robust control by tracking the geometric center position and orientation of objects with known shapes and sizes. However, this approach is limited in dynamic environments, especially when object shape, friction, and stiffness are unknown. Although friction is not explicitly modeled, we rely on posture-based closed loop control and low-frequency vibrations, which may be ineffective for high-friction (static coefficient around 0.8) or delicate objects. Future work will focus on integrating multi-modal perception to better estimate object properties and adopting learning-based control strategies that adapt to varying conditions and object types. We also plan to explore active friction modulation methods and low-friction materials for improved friction response. In terms of scalability, our current setup uses two robotic modules with 150×150 $mm^{2}$ top surfaces. We will investigate both upscaling and miniaturization—larger modules (up to 500×500 $mm^{2}$) for heavy-load tasks, requiring stronger actuators such as high-torque motors or hydraulics, and smaller modules (5×5–10×10 $mm^{2}$) for integration into robotic palms with actuation method and material selection adapted to the scale. Expanding to four or six modules will enlarge the workspace, enable more deformation modes, and support simultaneous manipulation of multiple objects, increasing efficiency for tasks like automated food handling. Finally, the dynamic manipulation principles demonstrated in prior works (e.g., tossing or flipping by momentum [R2cite2]) could also be integrated into our system, further extending its capabilities.
@@ 4.1 Surface-Based Integrated Design
The reconfigurable surface implemented in this project comprises two identical robotic modules. Each module has five main components (see Supplementary Fig. 1a). The surface platform is manufactured from an acrylic sheet, serving as the interface for contacting and manipulating objects via reconfigurability. The origami legs consist of a polyimide layer sandwiched between two FR4 layers, adhered together using two adhesive layers activated through a heat press application. The origami design derives from the fundamental waterbomb origami fold that can enable motion across three DoF, similar to the structures and their functionalities in the previous works [Salerno2020, mintchev2019portable, Mete2021, robertson2021soft, baines2023multi]. The legs facilitate the system’s reconfigurability, with each leg actuated by direct-driven BLDC motors (maxon EC-i 40) linked with 3D-printed PLA connectors. Finally, the base platform fabricated from an acrylic sheet secures the module onto the surface where the modules are fixed. Each motor, allocated to actuate the robotic modules, operates with a $36V$ nominal voltage and can generate a nominal torque of up to $207mN.m$. The STSPIN32F0A module (by STMicroelectronics) performs low-level control of the motors and manages the required current flow from the power supply. A main controller unit (Teensy MicroMod by PJRC and SparkFun) functions to handle high-level control operations and coordinates communication between these two boards. This communication utilizes the MODBUS RS-485 protocol and the Teensy controller’s TTL output. MAX485 module converts the main controller’s TTL output to the MODBUS RS-485 protocol, creating a coherent and reliable communication channel between the low-level actuation control and high-level functions.